OpenStudy (anonymous):

PLEASE PLEAS HELP I WILL MEDAL!!!! AND FAN how do you find the surface area and latreal of a clylinder?????

OpenStudy (anonymous):

area = 2 · π · r · (r + h)=surface area

OpenStudy (anonymous):

lateral area = 2*3.14*rh

OpenStudy (anonymous):

A=2πrh+2πr2

OpenStudy (anonymous):

A=2πrh+2πr^2

OpenStudy (anonymous):

I must be doing it wrong because ive tried both ways and i still get it wrong

OpenStudy (anonymous):

what are the figures?

OpenStudy (anonymous):

@lunasee

OpenStudy (anonymous):

A=2πrh (for lateral area)

OpenStudy (anonymous):

it says use the net to find the latreal area of the clyinder diamter is 12 hieght is 15

OpenStudy (anonymous):

also says give your answer in pi , which i dont understand

OpenStudy (anonymous):

im using the correct formula but i still get it wrong

OpenStudy (anonymous):

since the diameter is 12 the radius will be 6 since tha radius is half of diameter

OpenStudy (anonymous):

the answer i got for surface area is 791.68

OpenStudy (anonymous):

@lunasee

OpenStudy (anonymous):

thats the same that i got , it still says incorrect

OpenStudy (anonymous):

probably approximate it

OpenStudy (anonymous):

565.49 for the lateral area

OpenStudy (anonymous):

nope , neither work , ive tried it all it still says its incorrect , maybe because it says it needs to be in pi?

OpenStudy (anonymous):

@kalahkid

OpenStudy (anonymous):

i guessing you should put pi at the end of the answer??

OpenStudy (anonymous):

I have tried that too!!! This is driving me insane!!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

did you put the units cm or m^2 since its area

OpenStudy (anonymous):

@lunasee

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

Do you have to show working??

OpenStudy (triciaal):